OpenStudy (anonymous):

What is the y-intercept of the line? y=(1/3)x+2 and some of my answers are -2, -1/3, 1/3, 2 and I will give a medal

OpenStudy (noseboy908):

firstly, you need to know slope-intercept form. Do you have that with you?

OpenStudy (anonymous):

No noseboy908

OpenStudy (noseboy908):

ok, no problem. it's "y= mx+b", where m is the slope, and "b" is the y-intercept. so, what's the "b" in your equation?

OpenStudy (anonymous):

y = 1/3(x + 2) do you know how to distribute through the parenthesis ?

OpenStudy (anonymous):

I don't understand this at all

OpenStudy (anonymous):

distributing through the parenthesis....all you do is multiply the number outside of the parenthesis with everything in the parenthesis. y = 1/3(x + 2) basically, this is what you are doing.. y = 1/3*x + 1/3 * 2 I just multiplied 1/3(the number outside, I should say, right before the parenthesis, by everything in the parenthesis. So now we get .. y = 1/3x + 2/3 In y = mx + b form, the y intercept is in the b position...therefore, the y intercept is 2/3. Understand ??

OpenStudy (anonymous):

I don't have 2/3 for an answer

OpenStudy (anonymous):

oh...wait...sorry...I read the problem wrong... y = 1/3x + 2 In y = mx + b form, the y intercept is in the b position....so your y intercept is 2

OpenStudy (anonymous):

I realize that...don't know why I put the parenthesis where I did.

OpenStudy (anonymous):

Its okay

OpenStudy (anonymous):

its just been one of those days...lol

OpenStudy (anonymous):

I know what u mean same here

OpenStudy (anonymous):

Its gonna get better..think positive :)